find the value of sin(theta) in standard position if a point with coordinates
(-3, 2) lies on its terminal side.

Answer by RAY100(1637)

a rough sketch helps. Terminus is in quadrant 2 at (-3,2)
.
with x =-3,,and y=2, Hyp = sqrt(-3^2+2^2) = 3.6055
.
ref angle is tan^-1=2/3,,,,,theta = 33.69 degrees,,,,180-33.69 = 146.31 degrees,,,answer
.
sin 33.69 = .5547,,,,,sin 146.31 = .5547
.
check,,,sin theta = 2/3.6055=.5547,,,ok
,
theta = 33.69 and base angle is 146.31 degrees
.